A magazine company had a profit of $98,000 per year when it had 32,000 subscribers. When it obtained 38,000 subscribers, it had a profit of $126,500. Assume that the profit P is a linear function of the number of subscribers s.
(a) Find the function P.
P(s) = (i got 6.125s - 98000) but it said it was wrong

(b) What will the profit be if the company obtains 50,000 subscribers?
$
i also got 208250 and it said that was wrong too
(c) What is the number of subscribers needed to break even?

and once agin i got 16000 subscribers and got that wrong as well please help.
...........................................................................

The data in the table below show the distance, in feet, a ball travels for various swing speeds, in miles per hour, of a bat.
Bat speed (mph) Distance (ft)
40 201
45 212
50 239
60 274
70 300
75 326
80 332
(a) Find the linear regression equation for these data. (Round your coefficients to three decimal places. A graphing calculator is recommended.)
y =
(b) Using the regression model, what is the expected distance a ball will travel when the bat speed is 58 miles per hour? (Round your answer to the nearest foot.)
ft
this one im absolutly stuck on and keep getting everything wrong....

my email is crystalrainstone@yahoo.com

Answer by ewatrrr(24785)

$98,000 per year when it had 32,000 subscribers. When it obtained 38,000 subscribers, it had a profit of $126,500.
38,000 , 126,000
32,000, 98,000 m = 28,000/6,000 = 28/6 = 14/3
y -98000 = (14/3)(x - 32,000)
y = (14/3)x - 51,333.33
f(50) = $182, 000
..........
Using Excel Scatter Plot
.
a) y = 3.992x + 41.041
b) F(58) = 240.64 = 241 ft